Problem 10-3A On January 1, 2017, Evers Company purchased the following two machines for use in its production process Machine A: The cash price of this machine was $46,500. Related expenditures included: sales tax $3,700, shipping costs $100, insurance during shipping $50, installation and testing costs $70, and $200 of oil and lubricants to be used with the machinery during its first year of operations. Evers estimates that the useful life of the machine is 5 years with a $4,300 salvage value remaining at the end of that time period. Assume that the straight-line method of depreciation is used Machine B: The recorded cost of this machine was $180,000. Evers estimates that the useful life of the machine is 4 years with a $10,150 salvage value remaining at the end of that time period. Prepare te folowing for Machine A. (Round answers to 0 decimal places, e.g·s, 12s. credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts 1. The journal entry to record its purchase on January 1, 2017 2. The journal entry to record annual depreciation at December 31, 2017 No. Account Titles and Explanation Debit 2.
media%2Fcab%2Fcab451c9-08aa-47a8-ad6c-da
Which method used to calculate depreciation on Machine B reports the highest amount of depreciation expense in year 1 (2017)? The highest amount in year 4 (2020) The highest total amount over the 4-year period? Click if you would like to Show Work for this question: oen Show Work SIMILAR
Answer #1

Answer

  • Problem 10-3A

Working – MACHINE A

Cash price

$                46,500

Sales Tax

$                   3,700

Shipping Costs

$                      100

Insurance

$                         50

Installation and testing cost

$                         70

A

Total Cost of Machine A

$                50,420

B

Salvage value

$                   4,300

C = A - B

Depreciable base

$                46,120

D

Estimated life

                              5

E = C/D

Annual Depreciation

$                   9,224

--Entries asked for MACHINE A

Date

Accounts title

Debit

Credit

01-Jan-17

Equipment [Machine A]

$                50,420

   Cash

$               50,420

(Machine A purchased)

31-Dec-17

Depreciation Expense - Equipment

$                   9,224

   Accumulated Depreciation - Equipment

$                 9,224

(12 month depreciation)

  • Answers for MACHINE B

2017

2018

2019

2020

Straight Line

42463

42463

42463

42463

Declining balance

90000

45000

22500

12350

Units of Activity

66990 [43500 x $ 1.54]

49280 [32000 x 1.54]

31570 [20500 x 1.54]

21791 [ 14150 x 1.54]

Working

--Straight Line

A

Cost

$          180,000.00

B

Residual Value

$            10,150.00

C=A - B

Depreciable base

$          169,850.00

D

Life [in years]

4

E=C/D

Annual SLM depreciation

$            42,462.50

--Declining balance

A

Cost

$          180,000.00

B

Residual Value

$            10,150.00

C=A - B

Depreciable base

$          169,850.00

D

Life [in years]

4

E=C/D

Annual SLM depreciation

$            42,462.50

F=E/C

SLM Rate

25.00%

G=F x 2

DDB Rate

50.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

1

$           180,000.00

50.00%

$         90,000.00

$             90,000.00

2

$              90,000.00

50.00%

$         45,000.00

$             45,000.00

3

$              45,000.00

50.00%

$         22,500.00

$             22,500.00

4

$              22,500.00

$         12350.00

$            10150.00

--Unit sof activity

A

Cost

$          180,000.00

B

Residual Value

$            10,150.00

C=A - B

Depreciable base

$          169,850.00

D

Usage

                   110,150

E = C/D

Depreciation per …..

$                       1.54

  • Answers

--Highest depreciation in Year 1 = Declining balance $ 90000
--Highest amount in Year 4 = Straight Line
--Highest total amount over 4 year period = Straight Line